# Sina Rismanchian, Yasaman Razeghi, Sameer Singh, Shayan Doroudi

### v1

@@ 5.3 Multimodal Reasoning
The existing literature features a range of studies that evaluate these models using naturalistic images [20, 21, 2], yet humans naturally are able to reason over abstract shapes [8, 53, 42] and also many use cases of LMMs involve understanding abstract shapes and sketches [16, 35]. Moreover, unlike naturalistic images [32, 43], the relationship between language and abstract shapes is highly intertwined as minimal alterations in language can lead to different visual perceptions in humans [14, 26]. Overall, recent surveys on deep learning for mathematical reasoning [29, 44] have pointed out that most of the available datasets and benchmarks on multimodal reasoning often rely on visual question-answering frameworks. However, these methods fall short because they are usually trained on datasets composed of natural images, rather than on datasets tailored to the integration of vision and language for mathematical tasks.
The Multimodal Algorithmic Reasoning (MAR) task tests multi-modal models on fundamental skills understandable by children, focusing on interpreting visual and linguistic information to answer questions. Perhaps the most relevant work to ours is the paper by [7] in which they introduced a dataset with 101 multiple-choice questions inspired by the Math Kangaroo contest for 6 to 8-year-olds, involving images and texts that the model must analyze together. The task has been shown to be challenging for multimodal deep neural networks, and the following trials to solve the problem have gained less than 25% accuracy on the test set [51]. TurtleBench includes abstract geometric shapes, and the task only relies on knowledge and reasoning over a set of simple functions in the Python Turtle library. Our open-ended benchmark and its flexibility over different modalities make evaluating different aspects of multimodal reasoning in LMMs more reliable.
@@ 6 Discussion on Educational Implications
While LLMs and LMMs have sparked interest among education researchers in AI tools such as tutors [22] for students, our work cautions against using these models without thorough evaluation. Although students can learn turtle programming on various platforms without AI help (e.g., Code.org [11, 12]), the effectiveness of these models as tutors or copilots is uncertain due to current limitations. Our work suggests that educational researchers can engage in systematically benchmarking LMMs to ensure their reliability before integrating them into student learning processes.
@@ 7 Conclusions
This study introduces TurtleBench, the first of its kind in benchmarks that focus on converting visual inputs to code outputs. The evaluation results from TurtleBench reveal a significant disparity between how humans tackle turtle programming and how SOTA AI models perform in understanding simple geometric shapes, reasoning about these shapes, and converting such understandings into executable code [38]. This gap underscores the challenges that lie ahead in the quest to enhance AI’s comprehension and problem-solving abilities to match human levels. We believe that TurtleBench serves as a crucial tool in the evaluation of models, offering a clear benchmark testing the limits of LMMs.
@@ 8 Limitations
One of the limitations of our work is that we did not experiment with fine-tuning techniques to better understand multimodal reasoning abilities in these models and how they could be improved. However, we argue that our experiments demonstrate that poor performance in the models is perhaps not due to their unfamiliarity with the syntax of Turtle, but rather is more related to vision components and their reasoning abilities. We plan to experiment with fine-tuning techniques in future work with smaller models such as Llava-1.5-13B with newer architectures for vision towers [25] to examine the effectiveness of these techniques.
@@ A.1 Reasons of Failure
Shape identification error Shape Identification Error: The model fails to completely capture existent shapes in the input image, confusing or misattributing shapes 25%
Counting error Counting Error: The model inadequately counts the elements. 35%
Orientation error Orientation Error: The model fails to correctly determine the spatial relationships between different components of a shape 21%
Implementation error Implementation Error: The model’s generated code does not adhere to the pre-planned pseudocode, resulting in incorrect implementation. 45%
@@ A.2.1 Basic Prompt
[fontsize=, frame=single,linenos=false,breaklines,breaksymbol=,escapeinside=——,bgcolor=white]text In each task, the user provides an image of an abstract geometric shape or pattern and an instruction, you need to generate a code in Python Turtle that follows the user’s request.
@@
[fontsize=, frame=single,linenos=false,breaklines,breaksymbol=,escapeinside=——,bgcolor=white]text You are Turtle Geometrician, you are an expert in reasoning about images and generating code in Python Turtle using images You need to follow the steps below before generating the answer: (1) Describe the relevant information from the image needed to answer the question. List all relevant artifacts from the image. (2) Use the information described in (1) to reason about the problem by working step by step to arrive at the final piece of code. (3) Generate the final code. NEVER use ”pensize” function in your code.
@@ A.2.3 A Complete Example
[fontsize=, frame=single,linenos=false,breaklines,breaksymbol=,escapeinside=——,bgcolor=white]text System: You are Turtle Geometrician, you are an expert in reasoning about images and generating code in Python Turtle using images. You need to follow the steps below before generating the answer: (1) Describe the relevant information from the image needed to answer the question. List all relevant artifacts from the image. (2) Use the information described in (1) to reason about the problem by working step by step to arrive at the final piece of code. (3) Generate the final code. NEVER use ”pensize” function in your code.
@@ A.2.4 Arbitrary Output
[fontsize=, frame=single,linenos=false,breaklines,breaksymbol=,escapeinside=——,bgcolor=white]text You are an expert in reasoning about images and generating code in any language you prefer. You need to follow the steps below before generating the code that answers the user’s request: (1) Describe the relevant information from the image needed to answer the question. List all relevant information from the image. (2) Use the information described in (1) to reason about the problem by working step by step to arrive at the final piece of code. (3) Generate the final code. Your code can be in any visual language or library, such as Matplotlib, TikZ, etc.
@@ A.3.1 Prompt used
[fontsize=, frame=single,linenos=false,breaklines,breaksymbol=,escapeinside=——,bgcolor=white]text Suppose that I have a library named Rabbit in Python. Rabbit library has an object constructor named Rabbit which is an object that moves on the screen and draws lines. It only has these functions: aa(length): goes front or back (if the length is negative) and draws a line with the length of pixels. bb(degree): The rabbit turns its head right or left (if degree is negative). cc(radius, degree): creates an arc with the given radius for the given degree. If degree=360 it creates a circle. The center of the circle is in the left of the rabbit. pp(vanish): if vanish=True vanishes Rabbit object so if it moves does not draw anything, and if vanish=False, it appears the Rabbit object so if it moves draws on the screen. you call the functions on an object of Rabbit, such as r.aa(length) where r is an object of Rabbit. When r is created, it faces north (up) on the screen and it does not vanish, so it is in drawing mode.

### v2

@@ 5.3 Multimodal Reasoning
The existing literature features a range of studies that evaluate these models using naturalistic images (Jiang et al., 2022; Johnson et al., 2017; Antol et al., 2015), yet humans naturally are able to reason over abstract shapes (Chollet, 2019; Zhang et al., 2019; Spelke and Kinzler, 2007) and also many use cases of LMMs involve understanding abstract shapes and sketches (Forbus et al., 2011; Nie et al., 2020). Moreover, unlike naturalistic images (Marjieh et al., 2022; Sucholutsky and Griffiths, 2024), the relationship between language and abstract shapes is highly intertwined as minimal alterations in language can lead to different visual perceptions in humans (Dillon, 2023; Lin and Dillon, 2023). Overall, recent surveys on deep learning for mathematical reasoning Lu et al. (2022); Sun et al. (2023) have pointed out that most of the available datasets sand benchmarks on multimodal reasoning often rely on visual question-answering frameworks. However, these methods fall short because they are usually trained on datasets composed of natural images, rather than on datasets tailored to the integration of vision and language for mathematical tasks.
The Multimodal Algorithmic Reasoning (MAR) task tests multi-modal models on fundamental skills understandable by children, focusing on interpreting visual and linguistic information to answer questions. Perhaps the most relevant work to ours is the paper by Cherian et al. (2023) in which they introduced a dataset with 101 multiple-choice questions inspired by the Math Kangaroo contest for 6 to 8-year-olds, involving images and texts that the model must analyze together. The task has been shown to be challenging for multimodal deep neural networks, and the following trials to solve the problem have gained less than 25% accuracy on the test set Wu et al. (2023). TurtleBench includes abstract geometric shapes, and the task only relies on knowledge and reasoning over a set of simple functions in the Python Turtle library. Our open-ended benchmark and its flexibility over different modalities make evaluating different aspects of multimodal reasoning in LMMs more reliable.
@@ 6 Discussion on Educational Implications
While LLMs and LMMs have sparked interest among education researchers in AI tools such as tutors KhanAcademy (2024) for students, our work cautions against using these models without thorough evaluation. Although students can learn turtle programming on various platforms without AI help (e.g., Code.org (2024a, b)), the effectiveness of these models as tutors or copilots is uncertain due to current limitations. Our work suggests that educational researchers can engage in systematically benchmarking LMMs to ensure their reliability before integrating them into student learning processes.
@@ 7 Conclusions
This study introduces TurtleBench, the first of its kind in benchmarks that focus on converting visual inputs to code outputs. The evaluation results from TurtleBench reveal a significant disparity between how humans tackle turtle programming and how SOTA AI models perform in understanding simple geometric shapes, reasoning about these shapes, and converting such understandings into executable code Rismanchian et al. (2024). This gap underscores the challenges that lie ahead in the quest to enhance AI’s comprehension and problem-solving abilities to match human levels. We believe that TurtleBench serves as a crucial tool in the evaluation of models, offering a clear benchmark testing the limits of LMMs.
@@ 8 Limitations
One of the limitations of our work is that we did not experiment with fine-tuning techniques to better understand multimodal reasoning abilities in these models and how they could be improved. However, we argue that our experiments demonstrate that poor performance in the models is perhaps not due to their unfamiliarity with the syntax of Turtle, but rather is more related to vision components and their reasoning abilities. We plan to experiment with fine-tuning techniques in future work with smaller models such as Llava-1.5-13B with newer architectures for vision towers Li et al. (2024b) to examine the effectiveness of these techniques.
@@ A.1 Reasons of Failure
Shape identification error The model fails to completely capture existent shapes in the input image, confusing or misattributing shapes 25%
Counting error The model inadequately counts the elements. 35%
Orientation error The model fails to correctly determine the spatial relationships between different components of a shape 21%
Implementation error The model’s generated code does not adhere to the pre-planned pseudocode, resulting in incorrect implementation. 45%
@@ A.2.1 Basic Prompt
[fontsize=, frame=single,linenos=false,breaklines,breaksymbol=,escapeinside=||,bgcolor=white]text In each task, the user provides an image of an abstract geometric shape or pattern and an instruction, you need to generate a code in Python Turtle that follows the user’s request.
@@
[fontsize=, frame=single,linenos=false,breaklines,breaksymbol=,escapeinside=||,bgcolor=white]text You are Turtle Geometrician, you are an expert in reasoning about images and generating code in Python Turtle using images You need to follow the steps below before generating the answer: (1) Describe the relevant information from the image needed to answer the question. List all relevant artifacts from the image. (2) Use the information described in (1) to reason about the problem by working step by step to arrive at the final piece of code. (3) Generate the final code. NEVER use "pensize" function in your code.
@@ A.2.3 A Complete Example
[fontsize=, frame=single,linenos=false,breaklines,breaksymbol=,escapeinside=||,bgcolor=white]text System: You are Turtle Geometrician, you are an expert in reasoning about images and generating code in Python Turtle using images. You need to follow the steps below before generating the answer: (1) Describe the relevant information from the image needed to answer the question. List all relevant artifacts from the image. (2) Use the information described in (1) to reason about the problem by working step by step to arrive at the final piece of code. (3) Generate the final code. NEVER use "pensize" function in your code.
@@ A.2.4 Arbitrary Output
[fontsize=, frame=single,linenos=false,breaklines,breaksymbol=,escapeinside=||,bgcolor=white]text You are an expert in reasoning about images and generating code in any language you prefer. You need to follow the steps below before generating the code that answers the user’s request: (1) Describe the relevant information from the image needed to answer the question. List all relevant information from the image. (2) Use the information described in (1) to reason about the problem by working step by step to arrive at the final piece of code. (3) Generate the final code. Your code can be in any visual language or library, such as Matplotlib, TikZ, etc.
@@ A.3.1 Prompt used
[fontsize=, frame=single,linenos=false,breaklines,breaksymbol=,escapeinside=||,bgcolor=white]text Suppose that I have a library named Rabbit in Python. Rabbit library has an object constructor named Rabbit which is an object that moves on the screen and draws lines. It only has these functions: aa(length): goes front or back (if the length is negative) and draws a line with the length of pixels. bb(degree): The rabbit turns its head right or left (if degree is negative). cc(radius, degree): creates an arc with the given radius for the given degree. If degree=360 it creates a circle. The center of the circle is in the left of the rabbit. pp(vanish): if vanish=True vanishes Rabbit object so if it moves does not draw anything, and if vanish=False, it appears the Rabbit object so if it moves draws on the screen. you call the functions on an object of Rabbit, such as r.aa(length) where r is an object of Rabbit. When r is created, it faces north (up) on the screen and it does not vanish, so it is in drawing mode.